OpenStudy (anonymous):

A 5.20 bucket of water is accelerated upward by a cord of negligible mass whose breaking strength is 70.0 . If the bucket starts from rest, what is the minimum time required to raise the bucket a vertical distance of 12.5 without breaking the cord? please help..... @Mathematics

OpenStudy (anonymous):

5,20 kg, 70.0 N and 12,5 m. i have been trying for that for hours.

OpenStudy (anonymous):

Okay, well the force being exerted by the cord is m(a+g) if the bucket is being accelerated upward. if T = 70 N, then \[a = 70N / 5.2kg - 9.8m/s^2 = 3.66 m/s^2\] is the maximum acceleration that it could have. Starting from rest, \[\Delta y = 1/2at^2 \rightarrow t = \sqrt{2\Delta y/a} = \sqrt{2*12.5/3.66} = 2.61 s\]

OpenStudy (anonymous):

thanks a lot :)